Reduce to canonical form

uxx - uxy + uyy + ux = 0


1
Expert's answer
2021-05-24T18:13:08-0400
"A=1, B=-1, C=1"

"B^2-4AC=1-4=-3<0"

The given partial differential equation is elliptic.

 The characteristic equations are given by 


"\dfrac{dy}{dx}=\dfrac{B-\sqrt{B^2-4AC}}{2A}=\dfrac{-1-\sqrt{-3}}{2(1)}"

"=\dfrac{-1-i\sqrt{3}}{2}"

and



"\dfrac{dy}{dx}=\dfrac{B+\sqrt{B^2-4AC}}{2A}=\dfrac{-1+\sqrt{-3}}{2(1)}"

"=\dfrac{-1+i\sqrt{3}}{2}"


Integrating these two equations, we get 


"y=x(\dfrac{-1-i\sqrt{3}}{2})+c_1"

and


"y=x(\dfrac{-1+i\sqrt{3}}{2})+c_2"

Assume that


"\xi=y+x(\dfrac{1+i\sqrt{3}}{2})"

"\eta=y+x(\dfrac{1-i\sqrt{3}}{2})"

The transformation


"\alpha=\dfrac{1}{2}(\xi+\eta)"

"\beta=\dfrac{1}{2i}(\xi-\eta)"


"\alpha=y+x"

"\beta=\dfrac{\sqrt{3}}{2}x"

"\alpha_x=1, \alpha_y=1"

"\alpha_{xx}=0, \alpha_{xy}=0, \alpha_{yy}=0"

"\beta_x=\dfrac{\sqrt{3}}{2}, \beta_y=0"

"\beta_{xx}=0, \beta_{xy}=0, \beta_{yy}=0"

"A=1, B=-1, C=1,D=1, E=0, F=0, G=0"


"\bar{A}=A\alpha_x^2+B\alpha_x\alpha_y+C\alpha_y^2=1-1+1=1"

"\bar{B}=2A\alpha_x\beta_x+B(\alpha_x\beta_y+\alpha_y\beta_x)+2C\alpha_y\beta_y"

"=2(1)(\dfrac{\sqrt{3}}{2})-1(0+(1)\dfrac{\sqrt{3}}{2})+2(1)(0)=\dfrac{\sqrt{3}}{2}"


"\bar{C}=A\beta_x^2+B\beta_x\beta_y+2C\alpha_y\beta_y="

"=\dfrac{3}{4}+0+0=\dfrac{3}{4}"

"\bar{D}=A\alpha_{xx}+B\alpha_{xy}+C\alpha_{yy}+D\alpha_x+E\alpha_y=1"

"\bar{E}=A\beta_{xx}+B\beta_{xy}+C\beta_{yy}+D\beta_x+E\beta_y=\dfrac{\sqrt{3}}{2}"

"\bar{F}=F=0, \bar{G}=G=0"


"u_{\alpha\alpha}+\dfrac{\sqrt{3}}{2}u_{\alpha\beta}+\dfrac{\sqrt{3}}{2}u_{\beta\beta}+u_{\alpha}+\dfrac{\sqrt{3}}{2}u_\beta=0"